OpenStudy (anonymous):

how do i solve x^2-256=0 by factoring and applying the zero product property

OpenStudy (anonymous):

we know that \[a^2-b^2=(a-b)(a+b)\] using this formula write our question as \[x^2-16^2=0\] \[(x-16)(x+16)=0\] or \[x-16=0\] \[x+16=0\] \[x=16\] \[x=-16\]